### v2

Abstract
A nontrivial connected graph is matching covered if each edge belongs to some perfect matching. For most problems pertaining to perfect matchings, one may restrict attention to matching covered graphs; thus, there is extensive literature on them. A cornerstone of this theory is an ear decomposition result due to Lovász and Plummer. Their theorem is a fundamental problem-solving tool, and also yields interesting open problems; we discuss two such problems below, and we solve one of them.
A subgraph $H$ of a graph $G$ is conformal if GV(H)G-V(H) has a perfect matching. This notion is intrinsically related to the aforementioned ear decomposition theorem — which implies that each matching covered graph (apart from $K_{2}$ and even cycles) contains a conformal bisubdivision of $\theta$, or a conformal bisubdivision of $K_{4}$, possibly both. (Here, $\theta$ refers to the graph with two vertices joined by three edges.) This immediately leads to two problems: characterize $\theta$-free (likewise, $K_{4}$-free) matching covered graphs. A characterization of planar $K_{4}$-free matching covered graphs was obtained by Kothari and Murty [J. Graph Theory, 82 (1), 2016]; the nonplanar case is open.
We provide a characterization of $\theta$-free matching covered graphs that immediately implies a poly-time algorithm for the corresponding decision problem. Our characterization relies heavily on a seminal result due to Edmonds, Lovász and Pulleyblank [Combinatorica, 2, 1982] pertaining to the tight cut decomposition theory of matching covered graphs. As corollaries, we provide two upper bounds on the size of a $\theta$-free graph, namely, $m\leq 2n-1$ and $m\leq\frac{3n}{2}+b-1$, where $b$ denotes the number of bricks obtained in any tight cut decomposition of the graph; for each bound, we provide a characterization of the tight examples. The Petersen graph and $K_{4}$ play key roles in our results.
keywords:
perfect matchings, matching covered graphs, conformal minors, ear decompositions, tight cuts, theta graph
Dedicated to Professor Murty
1 Introduction and summary
All graphs considered in this paper are loopless; however, we allow multiple/parallel edges. For general graph theoretic terminology, we follow Bondy and Murty [1], and for terminology specific to matching theory, we follow Lucchesi and Murty [6]. This paper may be regarded as a sequel to Kothari and Murty [9]; however, it is self-contained. For a graph, we use $n$ for its order and $m$ for its size.
A graph is matchable if it has a perfect matching. A connected nontrivial graph is matching covered if each of its edges participates in some perfect matching. For instance, Schönberger (1927) proved that every 22-connected cubic graph is matching covered; Figure 1 shows all such graphs of order at most six. This may also be deduced using Tutte’s 11-factor Theorem stated below — where $\sf{c_{odd}}(H)$ denotes the number of odd components of a graph $H$.
(a) $\theta$
(b) $K_{4}$
(c) $C_{4}^{*}$
(d) $\overline{C_{6}}$
(e) $K_{3,3}$
Figure 1: all 22-connected cubic graphs of order six or less
Theorem $1.$
[Tutte’s Theorem]
A graph $G$ has a perfect matching if and only if $\sf{c_{odd}}(G-S)\leq|S|$ for each subset $S$ of $V(G)$.
For a matchable graph $G$, a subset $B$ of $V(G)$ is a barrier if $\sf{c_{odd}}(G-B)=|B|$. Using Tutte’s Theorem, one may easily prove the following.
Proposition 1.2.
For distinct vertices $u$ and $v$ of a matchable graph $G$, the graph GuvG-u-v is not matchable if and only if there exists a barrier containing both $u$ and $v$. ∎
This immediately yields the first part of the following result; the second part is easily proved.
Corollary 1.3.
[Characterization of Matching Covered Graphs]
A matchable graph is matching covered if and only if each of its barriers is stable; furthermore, if $B$ is a barrier of a matching covered graph $G$, then each component of GBG-B is odd. ∎
In fact, Kötzig (1957) proved the following fundamental result; see [15, 6].
Theorem 1.4.
[The Canonical Partition Theorem]
The maximal barriers of a matching covered graph partition its vertex set.
There is extensive literature on matching covered graphs; see [6]. Two important aspects of this theory are the ear decomposition theory and the tight cut decomposition theory, each of which plays an important role in our work; in the next section, we discuss the former that may be viewed as a refinement of Whitney’s ear decomposition theory for 22-connected graphs.
1.1 Ear Decomposition Theory
Let $H$ be a subgraph of a graph $G$. We say that $H$ is a conformal subgraph if GV(H)G-V(H) is matchable. The following is easily proved by considering the symmetric difference of two perfect matchings.
Proposition 1.5.
In a matching covered graph, any two adjacent edges participate in a conformal cycle. ∎
In fact, Little [13] proved the strengthening that any two edges of a matching covered graph participate in a conformal cycle.
Now, let $H$ be a subgraph of a graph $G$. A single ear of $H$ in $G$ is an odd path (in $G$) whose ends lie in $V(H)$ but is otherwise disjoint from $H$. For a bipartite graph $G$, a sequence of subgraphs $(G_{0},G_{1},\dots,G_{r})$ is said to be a bipartite ear decomposition of $G$ if (i) $G_{0}$ is isomorphic to $K_{2}$, (ii) $G_{i+1}=G_{i}\cup P_{i}$ where $P_{i}$ is a single ear of $G_{i}$ in $G$ for each $0\leq i<r$, and (iii) $G_{r}=G$. Figure 2 shows an example for the cube graph. It is easily verified that each subgraph in such a sequence (if it exists) is a conformal matching covered subgraph of $G$. Furthermore, Hetyei proved that every bipartite matching covered graph admits a bipartite ear decomposition; see [15, 6].
$P_{1}$$P_{0}$$P_{2}$$G_{0}$$P_{3}$$P_{4}$
Figure 2: a bipartite ear decomposition of the cube graph
Also, if $H$ is any conformal matching covered subgraph of a bipartite matching covered graph $G$, then there exists a bipartite ear decomposition of $G$ in which $H$ appears. However, in the case of the nonbipartite matching covered graphs, adding single ears is not sufficient to preserve the matching covered property. For example, if $G=K_{4}$ and $H=C_{4}$, then observe that adding any single ear (that is, an edge) to $H$ results in a subgraph that is not matching covered. This motivates the notion of a “double ear”.
For a subgraph $H$ of a graph $G$, a double ear of $H$ in $G$ is a pair of single ears of $H$ in $G$, say $R:=\{P,Q\}$, such that $P$ and $Q$ are vertex-disjoint. By an ear, we mean either a single ear or a double ear.
For a matching covered graph $G$, a sequence of matching covered subgraphs $(G_{0},G_{1},\dots,G_{r})$ is said to be an ear decomposition of $G$ if (i) $G_{0}$ is isomorphic to $K_{2}$, (ii) $G_{i+1}=G_{i}\cup R_{i}$ where $R_{i}$ is an ear of $G_{i}$ in $G$ for each $0\leq i<r$, and (iii) $G_{r}=G$. We say that $G_{i+1}$ is obtained from $G_{i}$ by adding the ear $R_{i}$. Figure 3 shows an ear decomposition of the Petersen graph; one may prove that, up to symmetry, the given ear decomposition is in fact unique. It is easily observed that each subgraph in such a sequence is indeed conformal. Lovász and Plummer [15] proved the following fundamental result; also see [6].
Theorem 1.6.
[Ear Decomposition Theorem]
Every matching covered graph admits an ear decomposition.
$G_{0}$$R_{0}$$R_{1}$$R_{3}$$R_{2}$
Figure 3: an ear decomposition of the Petersen graph
We remark that a double ear is added only when adding either of its constituent single ears results in a subgraph that is not matching covered. We now use this fact and the ear decomposition theory to make some observations. Before that, we need some terminology. Bisubdividing an edge refers to the operation of subdividing an edge by inserting an even number of subdivision vertices. A graph $H$ is a bisubdivision of a graph $J$ if it may be obtained from $J$ by bisubdividing any (possibly empty) subset of its edge set. The following is easily observed.
Proposition 1.7.
For a matching covered graph $J$, distinct from $K_{2}$, every bisubdivision of $J$ is matching covered. ∎
Let $(G_{0},G_{1},\dots,G_{r})$ denote an ear decomposition of a matching covered graph $G$. Observe that the second graph $G_{1}$ (if it exists) is obtained by adding a single ear to $K_{2}$, and is thus an even cycle. We now proceed to examine the third graph in any ear decomposition; assume $r\geq 2$. If $G_{2}=G_{1}\cup P$, where $P$ is a single ear of $G_{1}$, then it is easy to see that $G_{2}$ is a bisubdivision of $\theta$ — the graph shown in Figure 1(a). On the other hand, if $G_{2}=G_{1}\cup R$, where $R=\{P,Q\}$ is a double ear of $G_{1}$ then, using the fact that $G_{2}$ is matching covered, the reader may observe that $G_{2}$ is a bisubdivision of $K_{4}$. This proves the following.
Theorem 1.8.
$\theta-K_{4}$ Theorem]
Every matching covered graph, except $K_{2}$ and cycles, either has a conformal bisubdivision of $\theta$, or has a conformal bisubdivision of $K_{4}$, possibly both. ∎
The above theorem inspires the following technical definitions. A matching covered graph $J$ is a conformal minor of a matching covered graph $G$ if the latter has a conformal subgraph $H$ that is a bisubdivision of $J$; in this case, for the sake of brevity, we say that $G$ is $J$-based; otherwise, we say that $G$ is $J$-free. For instance, the triangular prism $\overline{C_{6}}$, shown in Figure 1(d), is $\theta$-based but $K_{4}$-free. On the other hand, the Petersen graph $\mathbb{P}$, shown in Figure 3, is $K_{4}$-based but $\theta$-free. The above theorem may be restated as follows: every matching covered graph, except $K_{2}$ and cycles, is either $\theta$-based or $K_{4}$-based, possibly both. This immediately leads us to the following problems.
Problem 1.
Characterize $\theta$-free matching covered graphs.
Problem 2.
Characterize $K_{4}$-free matching covered graphs.
In this paper, we solve Problem 1 by giving a structural characterization of $\theta$-free matching covered graphs. Our result provides an $\mathsf{NP}$-characterization of $\theta$-free graphs that leads to a poly-time algorithm, and thus implies that the corresponding decision problem lies in $\mathsf{P}$.
Note that, in an ear decomposition of a bipartite matching covered graph, a single ear is added at each step. Thus, the characterization of $\theta$-free bipartite graphs immediately follows from our discussion in the paragraph preceding Theorem 1.8.
Proposition 1.9.
[Bipartite $\theta$-free Graphs]
The only $\theta$-free bipartite matching covered graphs are $K_{2}$ and the even cycles. ∎
In fact, in the case of bipartite graphs, the following stronger statement is easily proved by using Proposition 1.5 and considering the symmetric difference of two appropriate perfect matchings; its proof is also implicitly contained in the proof of [15, Theorem 4.1.6].
Proposition 1.10.
In a bipartite matching covered graph, any three pairwise adjacent edges participate in a conformal bisubdivision of $\theta$. ∎
We also note the following easy consequence of Proposition 1.5.
Corollary 1.11.
The only non-simple $\theta$-free matching covered graph is the cycle $C_{2}$.∎
Interestingly, but perhaps not surprisingly, the Petersen graph plays a crucial role in the solution to Problem 1; let us argue that it is indeed $\theta$-free.
Proposition 1.12.
The Petersen graph $\mathbb{P}$ is $\theta$-free, and adding any edge to it results in a $\theta$-based graph.
Proof.
Clearly, one must exploit the symmetries of the Petersen graph $\mathbb{P}$. Since $\mathbb{P}$ has girth five and is non-hamiltonian, its only even cycles are 66-cycles and 88-cycles. If $C$ is any 88-cycle, then $\mathbb{P}-V(C)$ is $K_{2}$; whereas if $C$ is any 66-cycle, then $\mathbb{P}-V(C)$ is $K_{1,3}$. Thus, the conformal cycles of $\mathbb{P}$ are precisely its 88-cycles; consequently, every conformal bisubdivision of $\theta$ (if one exists) must include an 88-cycle as a subgraph. Figure 3 shows an 88-cycle of $\mathbb{P}$, say $C$, in orange and red. The reader may observe that if $Q$ is any single ear of $C$ in $\mathbb{P}$, then $C+Q$ contains an odd cycle. This proves that $\mathbb{P}$ is indeed $\theta$-free.
Now, let $u$ and $v$ denote distinct vertices of $\mathbb{P}$. If $u$ and $v$ are adjacent, by Corollary 1.11, $\mathbb{P}+uv$ is $\theta$-based. Now suppose that $u$ and $v$ are nonadjacent; by exploiting the symmetries of $\mathbb{P}$, it suffices to consider one such pair. Observe that, in $\mathbb{P}$, there are two internally-disjoint odd uvuv-paths whose union is a conformal 88-cycle $C$; thus $C+uv$ is a conformal bisubdivision of $\theta$ in $\mathbb{P}+uv$. This proves the second statement. ∎
Our solution to the nonbipartite case of Problem 1 requires the well-known tight cut decomposition theory of matching covered graphs that we describe in the next section. However, before that, we provide more context to the reader by briefly discussing other related problems and results. Theorem 1.8 is reminiscent of the following result of Lovász [14], which is significantly harder to prove.
Theorem 1.13.
$K_{4}-\overline{C_{6}}$ Theorem]
Every nonbipartite matching covered graph is either $K_{4}$-based, or $\overline{C_{6}}$-based, possibly both.
As noted earlier, the triangular prism $\overline{C_{6}}$, shown in Figure 1(d), is $\theta$-based; thus, Theorem 1.13 implies Theorem 1.8 in the case of nonbipartite graphs. Also, Theorem 1.13 leads to Problem 2 and the following.
Problem 3.
Characterize $\overline{C_{6}}$-free matching covered graphs.
Kothari and Murty [10, 9] solved Problems 2 and 3 for all planar graphs, and their work implies that the corresponding decision problems lie in $\mathsf{P}$; however, the nonplanar case of each is an open problem. In the next section, we shall briefly describe their approach and explain why the same does not work for Problem 1. Before that, we describe yet another similar problem.
The problem of characterizing Pfaffian graphs is perhaps one of the most important open problems in matching theory; we refer the reader to [15, 6, 16] for its significance and equivalent formulations. Kasteleyn [8] proved that all planar graphs are Pfaffian. On the other hand, Little [11] proved that a bipartite matching covered graph is Pfaffian if and only if it is $K_{3,3}$-free; this implies that the corresponding decision problem lies in $\mathsf{co}\text{-}\mathsf{NP}$. This leads one to the following problem.
Problem 4.
Characterize $K_{3,3}$-free matching covered graphs.
We remark that the above problem is equivalent to characterizing Pfaffian graphs only in the case of bipartite graphs; for instance, it can be shown that the Petersen graph is $K_{3,3}$-free as well as non-Pfaffian.
A complete characterization of $K_{3,3}$-free bipartite graphs was provided by Robertson, Seymour and Thomas [17], and independently by McCuaig [16], and their works imply a poly-time algorithm for deciding whether a bipartite graph is Pfaffian. Their characterizations also rely on the tight cut decomposition theory due to a result of Vazirani and Yannakakis [19], and independently of Little and Rendl [12].
We conclude this section with an innocent observation that distinguishes Problem 1 from all of the other problems discussed above.
Proposition 1.14.
Let $J$ be a simple matching covered graph. A matching covered graph $G$ is $J$-free if and only if its underlying simple graph is $J$-free. ∎
We remark that the assumption that $J$ is simple is crucial. For instance, the graphs shown in Figures 1(a) and 1(c) are $\theta$-based, but their underlying simple graphs are clearly $\theta$-free. We now proceed to discuss the aforementioned tight cut decomposition theory that is crucial to solve Problem 1 as well as the other such problems described above.
1.2 Tight Cut Decomposition Theory
For a subset $X$ of the vertex set of a graph $G$, we denote by $\partial(X)$ the cut associated with $X$ — that is, the set of edges that have one end in $X$ and the other end in $\overline{X}:=V(G)-X$. We refer to $X$ and $\overline{X}$ as the shores of the cut $C:=\partial(X)$, and we denote the graph obtained by shrinking the shore $\overline{X}$ to a single vertex $\overline{x}$ by $G/(\overline{X}\rightarrow\overline{x})$, or simply by $G/\overline{X}$. The two graphs $G/X$ and $G/\overline{X}$ are called the $C$-contractions of $G$. Figure 4 shows an example of a cut $C$, and the corresponding $C$-contractions.
For a vertex $v$, we simplify the notation to $\partial(v):=\partial(\{v\})$; such a cut is called trivial. Likewise, for a subgraph $H$, we simplify the notation to $\partial(H):=\partial(V(H))$. When $G$ is of even order, we say that a cut $C$ is odd if both shores have an odd number of vertices; otherwise, we say $C$ is even.
CCDD
Figure 4: a tight cut $C$ and $C$-contractions
1.2.1 Tight Cuts
A cut $C:=\partial(X)$ of a matching covered graph $G$ is a tight cut if $|C\cap M|=1$ for each perfect matching $M$. Observe that, if $C$ is a nontrivial tight cut, then $G/X$ and $G/\overline{X}$ are matching covered graphs of smaller order. If either of them has a nontrivial tight cut, then we can obtain two smaller graphs in the same manner. We perform this process repeatedly until we obtain a list of matching covered graphs — each of which is free of nontrivial tight cuts. This process is called the tight cut decomposition procedure. For instance, the cuts $C$ and $D$ shown in Figure 4 are tight, and the reader may verify that every application of the tight cut decomposition procedure (to the given graph) yields three copies of $K_{4}$ (up to multiple edges).
A matching covered graph that is free of nontrivial tight cuts is called a brace if it is bipartite, or a brick if it is nonbipartite. For instance, all of the cubic graphs shown in Figures 1, 2 and 3 are bricks and braces. In general, a matching covered graph may admit several applications of the tight cut decomposition procedure. However, Lovász proved the following remarkable result.
Theorem 1.15.
[Unique Tight Cut Decomposition Theorem]
Any two applications of the tight cut decomposition procedure to a matching covered graph $G$ yield the same list of bricks and braces, except for the multiplicities of edges.
In light of the above, for a matching covered graph $G$, by bricks and braces of $G$, we mean the underlying simple graphs produced by any application of the tight cut decomposition procedure. In particular, the above result implies that the number of bricks of $G$ is an invariant that is denoted by $b(G)$. This invariant plays a key role in matching theory; see [6]. For instance, for the graph $G$ shown in Figure 4, $b(G)=3$. It is worth noting that $G$ is bipartite if and only if $b(G)=0$.
We now proceed to state a characterization of bricks due to Edmonds, Lovász and Pulleyblank [7], which also implies (as per our discussion in the next Section 1.2.2) that the bricks and braces of a matching covered graph may be computed in poly-time. A matchable graph $G$ is bicritical if GuvG-u-v has a perfect matching for each pair $u,v\in V(G)$. A barrier is trivial if it comprises at most one vertex; otherwise it is nontrivial. Proposition 1.2 implies the following characterization.
Corollary 1.16.
[Characterization of Bicritical Graphs]
A matchable graph is bicritical if and only if each of its barriers is trivial. ∎
We are now ready to state the aforementioned characterization of bricks.
Theorem 1.17.
[Characterization of Bricks]
Bricks are precisely the 33-connected bicritical graphs of order four or more.
In light of the above result, most authors including [9] require bicritical graphs to have at least four vertices. Note that, as per our relaxed definition, $K_{2}$ (up to multiple edges) is the only bicritical graph that is bipartite, and of course, the only one that has precisely two vertices. It is easily proved that every bicritical cubic graph is in fact 33-connected; consequently, by the above theorem, each such graph, distinct from $\theta$, is a cubic brick. We remark that a matching covered graph of order four or more is bicritical if and only if its number of braces is zero.
In order to tackle Problems 2 and 3, Kothari and Murty first reduced them to the case of bricks by proving the following result.
Theorem 1.18.
Let $C$ denote a tight cut of a matching covered graph $G$, and let $J$ denote a cubic brick. Then, $G$ is $J$-free if and only if both $C$-contractions of $G$ are $J$-free.
Since cubic bricks are simple, combining the above with Proposition 1.14, we conclude the following.
Corollary 1.19.
For a cubic brick $J$, a matching covered graph $G$ is $J$-free if and only if each of its bricks is $J$-free. ∎
Since $K_{4}$ and $\overline{C_{6}}$ are cubic bricks, and since the bricks of a matching covered graph may be computed in poly-time, the above result implies that in order to solve Problems 2 and 3, it suffices to characterize $K_{4}$-free and $\overline{C_{6}}$-free bricks, respectively.
CCSS
(a) $T_{6}$ — a $\theta$-free graph
(b) a $C$-contraction $L^{*}$
(c) a marked $S$-component $L^{\prime}$
Figure 5: $C$-contraction and marked $S$-component of $T_{6}$
Naturally, one is tempted to wonder if the above approach may be employed to solve Problem 1. Unfortunately, however, neither Proposition 1.14 (as discussed earlier) nor Theorem 1.18 holds for $J=\theta$. Consider the graph $T_{6}$ shown in Figure 5; the reader may easily verify that $T_{6}$ is $\theta$-free whereas each of its $C$-contractions is $K_{4}$ plus an edge, and thus $\theta$-based. We shall return to this example soon.
One of the issues, as we realized, is that Theorem 1.18 deals with arbitrary tight cuts. Luckily, a result of Edmonds, Lovász and Pulleyblank (abbreviated henceforth to ELP), that is equivalent to Theorem 1.17, proves the existence of special types of nontrivial tight cuts (whenever a nontrivial tight cut exists); as we shall see in the next section, their result comes to our rescue in our pursuit of solving Problem 1.
1.2.2 ELP Cuts
Recall Corollary 1.3, and let $B$ denote a barrier of a matching covered graph $G$. Observe that if $L$ is any (odd) component of GBG-B, then $\partial_{G}(L)$ is a tight cut; such a tight cut is called a barrier cut. For an illustration, see the graphs shown in Figure 6. Observe that the graph $H$, obtained from $G$ by shrinking each component of GBG-B to a single vertex, is a bipartite matching covered graph. This fact, coupled with Lovász’s Unique Tight Cut Decomposition Theorem (1.15), proves the following.
CCBB
(a) a $\theta$-free graph
$B$
(b) a $\theta$-based graph
Figure 6: barrier cuts
Proposition 1.20.
Let $B$ denote a barrier of a matching covered graph $G$, and let $\mathcal{L}$ denote the set of components of GBG-B. Then:
$b(G)=\sum_{L\in\mathcal{L}}b(G/\overline{V(L)})$
The following result is one of our steps towards solving Problem 1; the reader may find it instructive to use it to see why one of the graphs, shown in Figure 6, is $\theta$-based whereas the other one is $\theta$-free (but, interestingly, it has a bisubdivision of $\theta$ that is not conformal).
Theorem 1.21.
[Recursing along a Barrier]
Let $B$ denote a barrier of a matching covered graph $G$, and let $J$ denote any bicritical cubic graph. Then, $G$ is $J$-free if and only if, for each component $L$ of GBG-B, the graph $G/\overline{V(L)}$ is $J$-free.
A proof of the above theorem appears in Section 2.1. It is well-known and easily proved that, in a bipartite matching covered graph, every tight cut is a barrier cut. However, this is not true for nonbipartite graphs; consider the cut shown in Figure 7(a). We now proceed to discuss another special type of tight cut.
CCSSuuvv
(a) $T_{6}$ — a $\theta$-free graph
$S$
(b) a $\theta$-based graph
$S$
(c) a $\theta$-based graph
Figure 7: 22-separation
Let $S:=\{u,v\}$ denote a 22-vertex-cut, of a matching covered graph $G$, that is not a barrier; we refer to $S$ as a 22-separation. It follows from Tutte’s Theorem (1.1) that $\sf{c_{odd}}(G-S)=0$. The reader may observe that, for any partition of GSG-S into two disjoint subgraphs $G_{1}$ and $G_{2}$, the cuts $\partial_{G}(V(G_{1})\cup\{u\})$ and $\partial_{G}(V(G_{2})\cup\{u\})$ are nontrivial tight cuts; such tight cuts are called 22-separation cuts. See the graphs shown in Figure 7 for illustrations.
Let us recall our discussion pertaining to the $\theta$-free graph $T_{6}$, shown in Figure 5, whose both $C$-contractions are $\theta$-based; this seems counter-intuitive in light of Theorem 1.18 that is applicable to all bicritical cubic graphs except $\theta$. To work around this issue, we use the following concept from Bondy and Murty [1]. For a 22-vertex-cut $S:=\{u,v\}$ of a graph $G$, the graph $L^{\prime}:=G[V(L)\cup S]+uv$, where $L$ is a component of GSG-S, is called a marked $S$-component of $G$; the newly added edge joining $u$ and $v$ is called the marker edge of $L^{\prime}$. As shown in Figure 5, for $T_{6}$, both marked $S$-components are $K_{4}$ (without multiple edges) and thus $\theta$-free. We now make some other related observations.
Let $S:=\{u,v\}$ be a 22-separation of a matching covered graph $G$, let $L$ denote a component of GSG-S, let $L^{\prime}$ denote the corresponding marked $S$-component, and let $L^{*}$ denote the tight cut contraction $G/(\overline{V(L)\cup\{u\}}\rightarrow v)$; see Figure 5. We invite the reader to compare $L^{\prime}$ and $L^{*}$, and observe that they are isomorphic except for the multiplicities of the edges joining $u$ and $v$ — each of which is at least one. This, coupled with Proposition 1.5, proves the following.
Proposition 1.22.
Let $S:=\{u,v\}$ denote a 22-separation of a matching covered graph $G$, let $L$ be a component of GSG-S, and let $L^{\prime}$ be the corresponding marked $S$-component. Then: (i) $L^{\prime}$ is matching covered, (ii) $L$ is matchable, and (iii) there exists a conformal (odd) uvuv-path in $L$. ∎
In light of Proposition 1.14, if $J$ is any simple matching covered graph, then $L^{\prime}$ is $J$-free if and only if $L^{*}$ is $J$-free. However, since our primary interest, namely $\theta$, is not simple, we must resort to marked $S$-components (instead of tight cut contractions) in the case of 22-separation cuts. Also, the above observations, coupled with Lovász’s Unique Tight Cut Decomposition Theorem (1.15), imply the following.
Proposition 1.23.
Let $S$ denote a 22-separation of a matching covered graph $G$, and let $\mathcal{L^{\prime}}$ denote the set of marked $S$-components of $G$. Then:
$b(G)=\sum_{L^{\prime}\in\mathcal{L^{\prime}}}b(L^{\prime})$
The following two results comprise our second step towards solving Problem 1. The first of these follows immediately from Proposition 1.22 (ii) and (iii); Figure 7(c) illustrates an example.
Proposition 1.24.
Let $S$ denote a 22-separation of a matching covered graph $G$. If GSG-S has three or more components, then $G$ is $\theta$-based. ∎
In light of the above statement, the technical assumption in the next statement is quite harmless, and in fact necessary. Its proof appears in Section $2.$
Theorem 1.25.
[Recursing across a 22-separation]
Let $S$ denote a 22-separation of a matching covered graph $G$, and let $J$ denote any 33-connected cubic graph. Assume that if $J=\theta$ then GSG-S has precisely two components. Then, $G$ is $J$-free if and only if each marked $S$-component of $G$ is $J$-free.
The reader may find it instructive to use the above result to explain why $T_{6}$ (shown in Figure 7(a)) is $\theta$-free, whereas the graph shown in Figure 7(b) is $\theta$-based. We are now ready to state the seminal result of ELP whose equivalence to Theorem 1.17 may be verified by the reader.
Theorem 1.26.
[ELP Theorem]
Every matching covered graph, that has a nontrivial tight cut, either has a nontrivial barrier cut or a 22-separation cut, possibly both.
In light of the above, by an ELP cut, we mean a nontrivial tight cut that is either a barrier cut or a 22-separation cut. It is worth noting that a matching covered graph may have a nontrivial tight cut that is not an ELP cut. For instance, the reader may verify that the tight cut $C$, shown in Figure 4, is not an ELP cut; however, the cut $D$ is an ELP cut. The reader may find it surprising that the original proof of the above theorem relied on linear programming duality. Szigeti [18] gave the first graph-theoretical proof. More recently, Carvalho, Lucchesi and Murty [5] gave yet another proof that relies heavily on the theory developed by them. As alluded to earlier, the tight cut decomposition of a matching covered graph is poly-time computable. This is precisely due to the Canonical Partition Theorem (1.4) and the ELP Theorem (1.26).
We now invite the reader to observe that Theorems 1.21 and 1.25, along with the ELP Theorem (1.26), imply Corollary 1.19 of Kothari and Murty. Furthermore, if $J$ is any bicritical cubic graph, then along with Proposition 1.24 and Theorem 1.9 (in the case of $J=\theta$), it follows that in order to characterize $J$-free matching covered graphs, it suffices to characterize $J$-free bricks. In particular, to solve Problem 1, it only remains to characterize the $\theta$-free bricks. This brings us to our final step towards solving Problem 1.
Theorem 1.27.
[Characterization of $\theta$-free Bricks]
The only $\theta$-free bricks are $K_{4}$ and the Petersen graph $\mathbb{P}$.
We prove the above result using the brick generation theorem of Carvalho, Lucchesi and Murty [4] in Section 2.3. On a related note, we state the following consequence of Proposition 1.9.
Corollary 1.28.
[Characterization of $\theta$-free Braces]
The only $\theta$-free braces are $K_{2}$ and $C_{2}$. ∎
Finally, we combine the above with Theorems 1.27, 1.26, 1.21 and 1.25, and Proposition 1.24, to obtain the following characterization of $\theta$-free graphs.
Theorem 1.29.
[Characterization of $\theta$-free Graphs]
A matching covered graph $G$ is $\theta$-free if and only if at least one of the following holds:
  • (i)
    either $G$ is $K_{2},C_{2},K_{4}$ or the Petersen graph $\mathbb{P}$, or
  • (ii)
    $G$ has a nontrivial barrier and for any such barrier, say $B$, the graph $G/\overline{V(L)}$ is $\theta$-free for each component $L$ of GBG-B, or
  • (iii)
    $G$ has a 22-separation and for any such 22-separation, say $S$, the graph GSG-S has precisely two components, and both marked $S$-components of $G$ are $\theta$-free. ∎
Our above theorem implies a poly-time algorithm to decide whether a given matching covered graph $G$ is $\theta$-free; we describe this briefly. If $G$ is a brick or a brace, we simply use statement (i). Otherwise, we compute the canonical partition of $G$; see Theorem 1.4. If $G$ has a maximal barrier $B$ that is nontrivial, we recurse on the $|B|$ smaller graphs as defined in statement (ii). Otherwise, $G$ is bicritical; see Corollary 1.16. By the ELP Theorem (1.26), $G$ has a 22-separation cut. We compute any 22-vertex-cut $S$ of $G$; since $G$ is bicritical, $S$ is a 22-separation. As per statement (iii), if GSG-S has three or more components, we conclude that $G$ is $\theta$-based; otherwise, we recurse on the two marked $S$-components of $G$. We refer the reader to [6] for a discussion on the facts that the (a) canonical partition may be computed in poly-time and (b) checking whether a graph is a brick or a brace is poly-time. We leave the runtime analysis to the reader.
Intuitively, one would expect an upper bound on the size of a $\theta$-free matching covered graph. This turns out to be true, and an easy consequence of our above characterization. We first define an operation. For two graphs $G_{1}$ and $G_{2}$ such that $G_{1}\cap G_{2}$ is precisely $K_{2}$, the graph $G$ obtained from $G_{1}\cup G_{2}$ by deleting the only edge of $G_{1}\cap G_{2}$ is called the $K_{2}$-sum of $G_{1}$ and $G_{2}$. We now use this operation to define two families of bicritical graphs; one of them, denoted by $\mathcal{T}_{0}$, is a proper subset of the other, denoted by $\mathcal{T}$.
The brace $C_{2}$, the bricks $K_{4}$ and the Petersen graph $\mathbb{P}$ belong to $\mathcal{T}$, and any graph obtained from two members of $\mathcal{T}$ using the $K_{2}$-sum operation belongs to $\mathcal{T}$. We now define $\mathcal{T}_{0}$. The brace $C_{2}$ and the brick $K_{4}$ belong to $\mathcal{T}_{0}$, and any graph obtained from two members of $\mathcal{T}_{0}$ using the $K_{2}$-sum operation belongs to $\mathcal{T}_{0}$. For instance, the graph $T_{6}$ shown in Figure 7(a) is the third smallest member of $\mathcal{T}_{0}$ as well as $\mathcal{T}$.
We are now ready to prove upper bounds on the size of $\theta$-free graphs along with the characterizations of corresponding tight examples.
Corollary 1.30.
Every $\theta$-free matching covered graph $G$ satisfies the following:
  1. (i)
    $m\leq\frac{3n}{2}+b-1$, and equality holds if and only if $G\in\mathcal{T}$,
  2. (ii)
    $b\leq\frac{n}{2}-1$, and equality holds if and only if $G\in\mathcal{T}_{0}$, and
  3. (iii)
    $m\leq 2n-2$, and equality holds if and only if $G\in\mathcal{T}_{0}$.
Proof.
We first invite the reader to observe that, since $\mathcal{T}_{0}\subset\mathcal{T}$, statement (iii) follows immediately from statements (i) and (ii).
We now prove statements (i) and (ii); in order to do so, we invoke Theorem 1.29 and proceed by induction on order. If statement (i) of Theorem 1.29 holds, the reader may verify that the desired conclusions hold.
Now suppose that statement (ii) of Theorem 1.29 holds, and let $B$ denote a nontrivial barrier. Observe that, by Corollary 1.16, $G$ is not bicritical; thus $G\notin\mathcal{T}$. For each $1\leq i\leq|B|$, let $L_{i}$ denote the components of GBG-B, and for each $i$, we let $L^{\prime}_{i}:=G/\overline{V(L_{i})}$, and let $n_{i},m_{i}$ and $b_{i}$ denote the order, size and the number of bricks, respectively, of $L^{\prime}_{i}$. Since each $L^{\prime}_{i}$ is a smaller $\theta$-free graph, by the induction hypothesis, $m_{i}\leq\frac{3n_{i}}{2}+b_{i}-1$ and $b_{i}\leq\frac{n_{i}}{2}-1$. Now, we use easy counting arguments and Proposition 1.20 to conclude that $G$ satisfies the claimed bounds with strict inequality:
$m=\sum_{i=1}^{|B|}m_{i}\leq\sum_{i=1}^{|B|}\left(\frac{3n_{i}}{2}+b_{i}-1\right)=\frac{3n}{2}+b-{|B|}<\frac{3n}{2}+b-1$
and
$b=\sum_{i=1}^{|B|}b_{i}\leq\sum_{i=1}^{|B|}\left(\frac{n_{i}}{2}-1\right)=\frac{n}{2}-|B|<\frac{n}{2}-1$
Finally, suppose that statement (iii) of Theorem 1.29 holds, and let $S$ denote a 22-separation. We let $L^{\prime}_{1}$ and $L^{\prime}_{2}$ denote the marked $S$-components of $G$; for each $i\in\{1,2\}$, let $n_{i},m_{i}$ and $b_{i}$ denote the order, size and the number of bricks, respectively, of $L^{\prime}_{i}$. Since both $L^{\prime}_{i}$s are smaller $\theta$-free graphs, by the induction hypothesis, $m_{i}\leq\frac{3n_{i}}{2}+b_{i}-1$, and equality holds if and only if $L^{\prime}_{i}\in\mathcal{T}$; furthermore, $b_{i}\leq\frac{n_{i}}{2}-1$, and equality holds if and only if $L^{\prime}_{i}\in\mathcal{T}_{0}$. Since each $L^{\prime}_{i}$ has order four or more, by Corollary 1.11, each of them is simple; consequently, $S$ is a stable set in $G$. Note that $m=m_{1}+m_{2}-2$ and $n=n_{1}+n_{2}-2$. By easy counting arguments and Proposition 1.23, we conclude that the bound in statement (i) holds:
$m\leq\left(\frac{3n_{1}}{2}+b_{1}-1\right)+\left(\frac{3n_{2}}{2}+b_{2}-1\right)-2=\frac{3n}{2}+3+b-4=\frac{3n}{2}+b-1$
Observe that equality holds above if and only if $m_{i}=\frac{3n_{i}}{2}+b_{i}-1$ for each $i\in\{1,2\}$. Consequently, equality holds if and only if $L^{\prime}_{1},L^{\prime}_{2}\in\mathcal{T}$, whence $G\in\mathcal{T}$.
Using the same arguments, we infer that the bound in statement (ii) holds:
$b=b_{1}+b_{2}\leq\left(\frac{n_{1}}{2}-1\right)+\left(\frac{n_{2}}{2}-1\right)=\frac{n+2}{2}-2=\frac{n}{2}-1$
Observe that equality holds above if and only if $b_{i}=\frac{n_{i}}{2}-1$ for each $i\in\{1,2\}$. Consequently, equality holds if and only if $L^{\prime}_{1},L^{\prime}_{2}\in\mathcal{T}_{0}$, whence $G\in\mathcal{T}_{0}$. This completes the proof of statements (i) and (ii), and thus of Corollary 1.30. ∎
We remark that the bound that appears in statement (ii) of the above corollary in fact holds for all matching covered graphs, and is easily proved. However, there are $\theta$-based matching covered graphs that also satisfy this bound with equality; one such example is $T_{6}+uv$ as per the labeling in Figure 7(a).
It follows immediately from statement (iii) of Corollary 1.30 that every $\theta$-free matching covered graph has a vertex of degree three or less. This immediately yields the following, where $\delta$ denotes the minimum degree.
Corollary 1.31.
Every matching covered graph, with $\delta\geq 4$, is $\theta$-based. ∎
We remark that the $K_{2}$-sum operation that appears in Corollary 1.30 is reminiscent of the $C_{4}$-sum (aka 44-cycle sum) operation that was crucial in the solution of Problem 4 for bipartite graphs — in the works of Robertson, Seymour and Thomas [17], and of McCuaig [16].
In the next section, we prove Theorems 1.21, 1.25 and 1.27, in that order.
2 Proofs
For a subdivision $H$ of a graph $J$ with $\delta(J)\geq 3$, we refer to the vertices of $H$ of degree three or more as its branch vertices, and to its remaining vertices as subdivision vertices. Note that each branch vertex corresponds to a unique vertex of $J$, and we use the same label to refer to both. In case $H$ is a bisubdivision of $J$, we may assign a parity to each edge of $H$ as described below.
One may associate each edge $uv\in E(J)$ with a unique odd path $P_{uv}$ in $H$, each of whose internal vertices (possibly none) is a subdivision vertex of $H$. We remark that this association is uniquely defined whenever $J$ is simple; however, since we are particularly interested in $J=\theta$, we may not assume its uniqueness. Now, let $P_{uv}:=w_{1}w_{2}\dots w_{2k}$ where $w_{1}:=u$ and $w_{2k}:=v$. We say that an edge $w_{i}w_{i+1}$ is of odd parity if $i$ is odd; otherwise of even parity. Note that the parity of an edge is independent of the order in which the path is traversed. In particular, the first and last edges (possibly not distinct) of any such path, say $P_{uv}$, are both of odd parity. Next, we assume $J$ to be matching covered, and relate perfect matchings of $H$ with those of $J$; recall Proposition 1.7.
Let $uv\in E(J)$. Note that, for every perfect matching $M_{J}$ of $J$ that contains uvuv, there is a corresponding perfect matching $M_{H}$ of $H$ that contains all edges of $P_{uv}$ of odd parity. Similarly, for each perfect matching $M_{J}$ of $J$ that does not contain uvuv, there is a corresponding perfect matching $M_{H}$ of $H$ that contains all edges of $P_{uv}$ of even parity. In fact, every perfect matching of $H$ arises in this manner, and $M_{J}\rightarrow M_{H}$ is a bijection between the sets of perfect matchings of $J$ and those of $H$.
Now, suppose that $H$ is a conformal subgraph of a matching covered graph $G$, and let $C$ denote any cut of $G$. For an edge uvuv of $J$, we say that the corresponding path $P_{uv}$ of $H$ is a $C$-crossing path if at least one edge of $P_{uv}$ lies in $C$, and we call such an edge a $C$-crossing edge. If there is one such edge, we say that $P_{uv}$ crosses $C$ once; if there are two such edges, we say that $P_{uv}$ crosses $C$ twice; and so on. We now state some observations due to Kothari and Murty [9]. They imposed stronger conditions on $J$; however, their proof goes through exactly as is with the weaker assumptions stated below.
Proposition 2.1.
Let $H$ denote a bisubdivision of a matching covered graph $J$, where $\delta(J)\geq 3$, that is conformal in a matching covered graph $G$, and let $C$ be a tight cut of $G$.
  • (i)
    For a $C$-crossing path $P_{uv}$, any two $C$-crossing edges of $P_{uv}$ must be of opposite parity. (Thus, $|C\cap~E(P_{uv})|\leq 2$).
  • (ii)
    If a $C$-crossing path $P_{uv}$ crosses $C$ twice, then there are no other $C$-crossing paths.
  • (iii)
    If $P_{uv}$ and $P_{uw}$ are two $C$-crossing paths, then each of them must cross $C$ in an edge of odd parity.
We now prove an easy consequence of the above that will be useful to us in the proof of the reverse implication of Theorem 1.21.
Corollary $2.$
Let $H$ be a bisubdivision of a matching covered graph $J$, where $\delta(J)\geq 3$, that is conformal in a matching covered graph $G$, and let $C:=\partial(X)$ be a nontrivial tight cut of $G$. If the shore $\overline{X}$ contains at most one branch vertex of $H$, then $G/\overline{X}$ is $J$-based.
Proof.
We let $G^{\prime}:=G/\overline{X}$, let $F:=\partial(X)\cap E(H)$, let $M_{H}$ denote a perfect matching of $H$, and $M$ be an extension of $M_{H}$ to a perfect matching of $G$. We consider a couple of cases and make some observations pertaining to the cardinalities of $M_{H}\cap C$ and $(M-M_{H})\cap C$.
First, suppose that there is a branch vertex $z$ in $\overline{X}$. Since $\delta(J)\geq 3$, by Proposition 2.1 (iii), each $C$-crossing path emanating from $z$ meets $C$ in an edge of odd parity. Consequently, $|M_{H}\cap C|=1$ and $|(M-M_{H})\cap C|=0$. Now, suppose that all branch vertices lie in $X$. By Proposition 2.1 (i) and (ii), $|F|\in\{0,2\}$. Furthermore, if $|F|=2$ then both edges belong to the same $C$-crossing path and are of opposite parities, and this implies that $|M_{H}\cap C|=1$ and $|(M-M_{H})\cap C|=0$. On the other hand, if $|F|=0$, then $|M_{H}\cap C|=0$ and $|(M-M_{H})\cap C|=1$.
In all cases discussed above, observe that the subgraph $H^{\prime}:=G^{\prime}\cap E(H)$ is a bisubdivision of $J$ in $G^{\prime}$, and $M^{\prime}:=M\cap E(G^{\prime})$ is a perfect matching of $G^{\prime}-V(H^{\prime})$. This completes the proof of Corollary $2.$ ∎
2.1 Dealing with barriers: a proof of Theorem 1.21
We shall first prove the forward implication of Theorem 1.21 with the weaker hypothesis that $J$ is a matching covered graph with maximum degree $\Delta(J)\leq 3$. Before doing so, we prove a couple of technical results.
Lemma 2.3.
Let $H$ denote a bisubdivision of a matching covered graph $J$, where $\delta(J)\geq 3$, that is conformal in a matching covered graph $G$, and let $B$ denote a barrier of $G$. If GBG-B has distinct components, say $K$ and $L$, each of which contains exactly one branch vertex of $H$, say $u$ and $v$, respectively, then $u$ and $v$ are nonadjacent in $J$.
Proof.
Suppose to the contrary that $u$ and $v$ are adjacent in $J$, and let $P_{uv}$ denote the path of $H$ corresponding to one edge joining $u$ and $v$ in $J$. Since $\delta(J)\geq 3$, there are three or more $\partial(K)$-crossing paths emanating from $u$; by Proposition 2.1 (iii), each of them crosses $\partial(K)$ in an edge of odd parity. In particular, $P_{uv}$ meets $\partial(K)$ in an edge of odd parity, say $e_{1}$. By an analogous argument, $P_{uv}$ meets $\partial(L)$ in an edge of odd parity, say $e_{2}$. Consequently, the subpath $e_{1}P_{uv}e_{2}$ is an odd path. We intend to arrive at a contradiction to this.
Note that if $P_{uv}$ meets a component $F$ of GBG-B that is distinct from $K$ and $L$, then by Proposition 2.1 (i), $P_{uv}$ has precisely two $\partial(F)$-crossing edges and they are of opposite parity; thus, the subpath $P_{uv}\cap F$ is even. Using this observation, and the fact that $B$ is a stable set, we conclude that the subpath $e_{1}P_{uv}e_{2}$ is an even path; this contradicts what we have already established above. ∎
We now use the above result to prove a generalization of Proposition 1.10.
Proposition 2.4.
Let $B$ denote a barrier of a matching covered graph $G$, and let $z$ denote an isolated vertex of GBG-B. Then, any three edges in $\partial(z)$ participate in a conformal bisubdivision of $\theta$.
Proof.
We proceed by induction on the order of $G$. Let $K$ denote a set of three edges incident at $z$. If each (odd) component of GBG-B is trivial, then $G$ is bipartite, and using Proposition 1.10 we observe that $K$ participates in a conformal bisubdivision of $\theta$.
Now let $L$ be a nontrivial odd component of GBG-B. Let $U:=V(L)$, let $L_{1}:=G/U\rightarrow u$, and let $L_{2}:=G/\overline{U}\rightarrow\overline{u}$. Observe that, in $L_{1}$, the set $B$ is a barrier, and $z$ is an isolated vertex of $L_{1}-B$. By the induction hypothesis, $K$ participates in a conformal bisubdivision of $\theta$, say $H_{1}$, in $L_{1}$; let $M_{1}$ be a perfect matching of $L_{1}-V(H_{1})$. By Lemma 2.3, the branch vertex of $H_{1}$, distinct from $z$, lies in $B$; in particular, $u$ is not a branch vertex of $H_{1}$.
Now, let $F:=E(H_{1})\cap\partial(U)$. It follows from the above, and Proposition 2.1 (i) and (ii), that $|F|\in\{0,2\}$; furthermore, if $|F|=2$, then the two edges belong to the same $\partial(U)$-crossing path of $H_{1}$, and are of opposite parities. If $|F|=0$, let $H_{2}$ denote the null subgraph of $L_{2}$ and let $M_{2}$ denote a perfect matching of $L_{2}$ that contains the unique edge in $M_{1}\cap\partial(U)$. On the other hand, if $|F|=2$, we invoke Proposition 1.5, and let $H_{2}$ denote a conformal cycle in $L_{2}$ containing $F$, and let $M_{2}$ denote a perfect matching of $L_{2}-V(H_{2})$. Observe that, in each case, $H:=H_{1}\cup H_{2}$ is a bisubdivision of $\theta$ in $G$, and $M:=M_{1}\cup M_{2}$ is a perfect matching of GV(H)G-V(H). ∎
We are now ready to prove the forward implication of Theorem 1.21.
Theorem 2.5.
Let $B$ denote a barrier of a matching covered graph $G$, and let $J$ be a matching covered graph with $\Delta(J)\leq 3$. If there exists an (odd) component $L$ of GBG-B such that $G/\overline{V(L)}$ is $J$-based, then $G$ is also $J$-based.
Proof.
Since every matching covered graph is $K_{2}$-based, we may assume that $J\neq K_{2}$; henceforth, $\delta(J)\geq 2$. Let $H_{1}$ denote a conformal bisubdivision of $J$ in $L_{1}:=G/(Z\rightarrow z)$, where $Z:=\overline{V(L)}$. Let $M_{1}$ denote a perfect matching of $L_{1}-V(H_{1})$. We let $L_{2}$ denote the other $\partial(Z)$-contraction; that is $L_{2}:=G/\overline{Z}$. Let $F:=E(H_{1})\cap\partial(Z)$. Since $2\leq\delta(J)\leq\Delta(J)\leq 3$, we infer that $|F|\in\{0,2,3\}$.
Depending on the cardinality of $F$, we will define a conformal subgraph $H_{2}$ in $L_{2}$, and a perfect matching $M_{2}$ of $L_{2}-V(H_{2})$. If $F=\emptyset$, we let $H_{2}$ be the null subgraph of $L_{2}$, and we let $M_{2}$ denote a perfect matching of $L_{2}$ that contains the unique edge in $M_{1}\cap\partial(Z)$. If $|F|=2$, we invoke Proposition 1.5, and let $H_{2}$ denote a conformal cycle in $L_{2}$ containing $F$, and let $M_{2}$ be any perfect matching of $L_{2}-V(H_{2})$. If $|F|=3$, we invoke Proposition 2.4, and let $H_{2}$ denote a conformal bisubdivision of $\theta$ containing $F$ in $L_{2}$, and let $M_{2}$ be any perfect matching of $L_{2}-V(H_{2})$. In all three cases, the reader may verify that $H:=H_{1}\cup H_{2}$ is a bisubdivision of $J$ in $G$, and that $M:=M_{1}\cup M_{2}$ is a perfect matching of GV(H)G-V(H). ∎
The reverse implication of Theorem 1.21 turns out to be more difficult to prove. We thus need a few more technical results; these, or weaker versions of them, were also used by Kothari and Murty [9]. The first of these is a special case of a result due to Plesnik [15, Theorem 3.4.2].
Lemma 2.6.
If $J$ is any 22-connected cubic graph, then $J-e_{1}-e_{2}$ is matchable for each pair of edges $e_{1}$ and $e_{2}$.
The next result is proved in [9] for all bricks, and it holds trivially for matching covered graphs of order two.
Lemma 2.7.
In a 33-connected bicritical graph $J$, for any nontrivial cut $C$ and edge $e\in C$, there exists a perfect matching $M_{J}$ such that $|M_{J}\cap(C-e)|\geq 2$.
We now prove a consequence of the above two lemmas. Recall from Section 1.2.2 that every bicritical cubic graph is 33-connected.
Corollary 2.8.
Let $H$ denote a bisubdivision of a bicritical cubic graph $J$ that is conformal in a matching covered graph $G$, and let $C$ be a nontrivial tight cut of $G$. If one shore of $C$ contains two or more branch vertices of $H$, then the other shore contains at most one branch vertex of $H$.
Proof.
Since $C$ is tight, and each perfect matching $M_{H}$ of $H$ extends to a perfect matching of $G$, we conclude that $|M_{H}\cap C|\leq 1$. In what follows, we shall invoke this observation frequently. We let $C:=\partial(X)$, and we let $Y$ and $\overline{Y}$ denote the sets of branch vertices of $H$ that are contained in $X$ and $\overline{X}$, respectively. Our goal is to prove that either $|Y|\leq 1$ or $|\overline{Y}|\leq 1$.
We begin by making an observation that is applicable when $Y$ and $\overline{Y}$ are both nonempty. Since $J$ is 33-connected as noted earlier, $H$ has at least three $C$-crossing paths; by Proposition 2.1 (i) and (ii), each of them crosses $C$ in exactly one edge. If two of these $C$-crossing edges, say $e_{1}$ and $e_{2}$, are of even parity, then Lemma 2.6 implies that there exists a perfect matching $M_{H}$ of $H$ that contains both $e_{1}$ and $e_{2}$, and thus $|M_{H}\cap C|\geq 2$, contrary to our first observation. Hence, we conclude that at most one $C$-crossing edge is of even parity. If there is such an edge $e$, we let $P_{uv}$ denote the $C$-crossing path that contains $e$ where $uv\in E(J)\cap\partial_{J}(Y)$. Otherwise, we let uvuv be any edge of $E(J)\cap\partial_{J}(Y)$.
Suppose to the contrary that $|Y|\geq 2$ and $|\overline{Y}|\geq 2$. By Lemma 2.7, there exists a perfect matching $M_{J}$ of $J$ satisfying $|M_{J}\cap(\partial_{J}(Y)-uv)|\geq 2$. Let $M_{H}$ denote the corresponding perfect matching of $H$. By our choice of uvuv, we infer that $|M_{H}\cap C|\geq 2$, contrary to our first observation. This completes the proof of Corollary 2.8. ∎
We are now ready to prove the reverse implication of Theorem 1.21.
Theorem 2.9.
Let $J$ denote a bicritical cubic graph, and let $B$ denote a barrier of a matching covered graph $G$. If $G$ is $J$-based then $G/\overline{V(L)}$ is $J$-based for some component $L$ of GBG-B.
Proof.
Let $H$ denote a conformal bisubdivision of $J$ in $G$. We shall first locate the branch vertices of $H$, and then invoke Corollary 2.2 to finish the proof.
2.9.1.
At most one branch vertex of $H$ lies in the barrier $B$.
Proof.
Since JuvJ-u-v has a perfect matching for each pair $u,v\in V(J)$, the reader may observe that HuvH-u-v has a perfect matching, and by conformality of $H$, the graph GuvG-u-v has a perfect matching. Consequently, by Proposition 1.2, at most one of $u$ and $v$ lies in $B$. ∎
2.9.2.
If there exist two or more components of GBG-B that contain branch vertices of $H$, then there exists a component of GBG-B that contains at least two branch vertices of $H$.
Proof.
Suppose to the contrary that at least two components of GBG-B contain branch vertices of $H$, but each such component contains exactly one branch vertex. Note that either $J=\theta$ or otherwise each vertex of $J$ has at least two neighbors. Thus, it follows from 2.9.1 that there exist components $K$ and $L$ of GBG-B that contain branch vertices $u$ and $v$, respectively, such that $u$ and $v$ are adjacent in $J$. This contradicts Lemma 2.3. ∎
We remark that if $J=\theta$, then the above statement implies that either both branch vertices belong to a particular component of GBG-B, or otherwise precisely one of them lies in $B$.
2.9.3.
There exists a component $L$ of GBG-B that contains at least |V(J)|1|V(J)|-1 branch vertices.
Proof.
The statement holds trivially for $J=\theta$. Otherwise, using 2.9.2, there exists a component $L$ of GBG-B that contains at least two branch vertices of $H$. Observe that $\partial(L)$ is a nontrivial tight cut; since $J$ is bicritical and cubic, we invoke Corollary 2.8 to infer that $\overline{V(L)}$ contains at most one branch vertex. ∎
Since $\partial(L)$ is a tight cut, we invoke Corollary 2.2 to conclude that $G/\overline{V(L)}$ is $J$-based, and this completes the proof of Theorem 2.9. ∎
The above theorem, along with Theorem 2.5, prove Theorem 1.21.
2.2 Handling 22-separations: a proof of Theorem 1.25
In this section, our goal is to prove Theorem 1.25. We first prove the forward implication with weaker assumptions.
Theorem 2.10.
Let $J$ and $G$ be matching covered graphs, and let $S$ denote a 22-separation of $G$. If some marked $S$-component of $G$ is $J$-based, then $G$ is also $J$-based.
Proof.
Let $S:=\{u,v\}$, and let $K$ and $L$ denote two distinct (even) components of GSG-S. Suppose the marked $S$-component $K^{\prime}$ (corresponding to $K$) is $J$-based, and let $H^{\prime}$ denote a conformal bisubdivision of $J$ in $K^{\prime}$. Let $e$ denote the marker edge of $K^{\prime}$, and let $P$ denote a conformal odd uvuv-path in $L$ provided by Proposition 1.22 (iii). Depending on whether or not $e$ belongs to $H^{\prime}$, either $H:=H^{\prime}+P-e$ or $H:=H^{\prime}$, respectively, is a conformal bisubdivision of $J$ in $G$. ∎
We now prove the reverse implication of Theorem 1.25 that turns out to be more difficult.
Theorem 2.11.
Let $S:=\{u,v\}$ denote a 22-separation of a matching covered graph $G$, and let $J$ denote any 33-connected cubic graph. Assume that if $J=\theta$ then GSG-S has precisely two components. If $G$ is $J$-based then some marked $S$-component of $G$ is also $J$-based.
Proof.
We let $H$ denote a bisubdivision of $J$ that is conformal in $G$. Since $H$ is a subdivision of a 33-connected graph, and since $S$ is a 22-separation of $G$, we immediately conclude that there do not exist two components of GSG-S each of which contains a branch vertex of $H$. This proves the following.
2.11.1.
There exists a marked $S$-component of $G$ that contains all branch vertices of $H$. ∎
Since $\theta$ is the only 33-connected cubic graph of order two, if $J\neq\theta$ then there exists a unique marked $S$-component of $G$, denoted by $L^{\prime}$, that contains all branch vertices of $H$. On the other hand, if $J=\theta$ then, by our hypothesis, there are precisely two marked $S$-components; in this case, each of them may contain both branch vertices, and this happens if and only if $u$ and $v$ are the branch vertices of $H$. So, if $J=\theta$, we let $L^{\prime}$ denote the marked $S$-component that minimizes the number of $\partial_{G}(V(L^{\prime}))$-crossing paths. We let $e$ denote the marker edge of $L^{\prime}$, let $D$ denote the even cut $\partial_{G}(V(L^{\prime}))$, and let $M$ denote a perfect matching of GV(H)G-V(H).
2.11.2.
$H$ has at most one $D$-crossing path.
Proof.
Suppose to the contrary that $H$ has at least two $D$-crossing paths, say $P$ and $Q$. Since all branch vertices belong to $V(L^{\prime})$, and since $S:=\{u,v\}$ is a 22-separation of $G$, each of $P$ and $Q$ contains $u$ as well as $v$. This immediately implies that $u$ and $v$ are branch vertices of $H$. Furthermore, since $J$ is 33-connected and cubic, we conclude that $J=\theta$. Consequently, $H=P\cup Q\cup R$, where $R$ is the third uvuv-path. Observe that if $K^{\prime}$ is the marked $S$-component of $G$ distinct from $L^{\prime}$, then the number of $\partial_{G}(V(K^{\prime}))$-crossing paths is at most one, and this contradicts our choice of $L^{\prime}$. ∎
We shall argue that $L^{\prime}$ is $J$-based. To this end, we consider a couple of cases; in each of them, we define a subgraph $H^{\prime}$ of $L^{\prime}$ and a matching $M^{\prime}$ of $L^{\prime}$.
First, suppose that $H$ has no $D$-crossing paths; we let $H^{\prime}:=H$. Observe that $|M\cap D|\in\{0,2\}$; if $|M\cap D|=0$, we let $M^{\prime}:=M\cap E(L^{\prime})$, and otherwise, we let $M^{\prime}:=(M\cap E(L^{\prime}))+e$. Now, suppose that $H$ has precisely one $D$-crossing path, say $P_{xy}$; clearly, $|M\cap D|=0$. We let $Q:=P_{xy}\cap(E(G)-E(L^{\prime}))$, let $H^{\prime}:=H-Q+e$ and $M^{\prime}:=M\cap E(L^{\prime})$.
In each case discussed above, the reader may verify that $H^{\prime}$ is a bisubdivision of $J$ and $M^{\prime}$ is a perfect matching of $L^{\prime}-V(H^{\prime})$. Thus, $L^{\prime}$ is $J$-based, and this completes the proof of Theorem 2.11. ∎
The above theorem, along with Theorem 2.10, proves Theorem 1.25.
2.3 Characterizing $\theta$-free bricks: a proof of Theorem 1.27
In this section, our goal is to characterize $\theta$-free bricks. To this end, we find the brick generation theorem due to Carvalho, Lucchesi and Murty [4] useful. In order to state their result, we need some terminology.
An edge $e$ of a matching covered graph $G$ is removable if GeG-e is also matching covered. In this case, if $G$ is a brick then GeG-e need not be a brick; in particular, GeG-e may have one or two vertices of degree two. In order to recover a brick, at the very least, we need to get rid of the degree two vertices (if any).
Let $v_{0}$ denote a vertex of degree two in a matching covered graph $G$ that has two distinct neighbors, say $v_{1}$ and $v_{2}$. Let $H$ be obtained from $G$ by contracting the two edges $v_{0}v_{1}$ and $v_{0}v_{2}$; we say that $H$ is obtained from $G$ by bicontracting the vertex $v_{0}$. Observe that $H=G/X$ where $X:=\{v_{0},v_{1},v_{2}\}$ and $\partial(X)$ is a barrier cut. The following is an immediate consequence of Theorem 1.21.
Corollary 2.12.
Let $G$ be a matching covered graph that has a vertex $v_{0}$ of degree two with two distinct neighbors, and let $H$ be obtained from $G$ by bicontracting the vertex $v_{0}$. Then, $G$ is $\theta$-free if and only if $H$ is $\theta$-free. ∎
$e$
Figure 8: bicorn and its thin edge $e$
Now, let $e$ be a removable edge of a brick $G$. As noted earlier, GeG-e has zero, one or two vertices of degree two. The graph $H$ obtained from GeG-e by bicontracting each of these degree two vertices is called the retract of GeG-e. Observe that $|V(G)|-|V(H)|\in\{0,2,4\}$; furthermore, each contraction vertex of $H$ (resulting from the bicontraction of degree two vertices), if any, has degree at least four. We say that $e$ is a thin edge of the brick $G$ if the retract of GeG-e is also a brick. For instance, the bicorn shown in Figure 8 has a unique removable edge that is also thin. On the other hand, if $e$ is any edge of the Petersen graph $\mathbb{P}$, the retract of $\mathbb{P}-e$ is precisely $T_{6}$; consequently, $\mathbb{P}$ is devoid of thin edges.
We are now ready to state the aforementioned generation theorem of Carvalho, Lucchesi and Murty [4] — that is a stronger version of a conjecture of Lovász proved earlier by the same authors [2, 3].
Theorem 2.13.
[Thin Edge Theorem]
Every brick distinct from $K_{4}$, $\overline{C_{6}}$ and the Petersen graph $\mathbb{P}$ has a thin edge.
In order to apply the above theorem, we will find the generation viewpoint useful. This viewpoint states that every brick may be generated from at least one of $K_{4},\overline{C_{6}}$ and $\mathbb{P}$ by means of a sequence of ‘expansion operations’. We now proceed to formalize this.
Let $v$ be a vertex of degree four or more in a matching covered graph $H$. Let $G$ be any graph obtained from $H$ by replacing $v$ with two new vertices $v_{1}$ and $v_{2}$, distributing the edges incident at $v$ amongst $v_{1}$ and $v_{2}$ so that each of them receives at least two edges, and introducing a new vertex $v_{0}$ and edges $v_{0}v_{1}$ and $v_{0}v_{2}$. We say that $G$ is obtained from $H$ by bisplitting the vertex $v$. Observe that $H$ may be recovered from $G$ by bicontracting the vertex $v_{0}$.
In their work [4], Carvalho, Lucchesi and Murty defined four expansion operations — each of which comprises a sequence of zero, one or two bisplittings followed by adding an edge. In particular, if $e$ is a thin edge of a brick $G$, and $H$ is the retract of GeG-e, then $G$ may be obtained from $H$ using one of the four expansion operations. In our work, we do not require the precise definitions of these operations; the interested reader may refer to [4]. We shall simply use the facts that the constituent operations of adding an edge as well as of bisplitting preserve the $\theta$-based property; the former is easy to see, whereas the latter follows from Corollary 2.12. We are now ready to prove our characterization of $\theta$-free bricks restated below.
Theorem 1.27.
[Characterization of $\theta$-free Bricks]
The only $\theta$-free bricks are $K_{4}$ and the Petersen graph $\mathbb{P}$.
Proof.
By Proposition 1.12, $\mathbb{P}$ is $\theta$-free. Clearly, $K_{4}$ is $\theta$-free. Conversely, let $G$ be a $\theta$-free brick. We proceed by induction on the size of $G$. If $G\in\{K_{4},\overline{C_{6}},\mathbb{P}\}$, since $\overline{C_{6}}$ is $\theta$-based, the desired conclusion holds. Otherwise, by the Thin Edge Theorem (2.13), $G$ has a thin edge, say $e$, and let $H$ denote the retract of GeG-e. Clearly, GeG-e is $\theta$-free; by Corollary 2.12, $H$ is also $\theta$-free. Since $H$ is a brick, by the induction hypothesis, $H\in\{K_{4},\mathbb{P}\}$. Furthermore, $G$ is obtained from $H$ by one of the four expansion operations. Since $H$ is cubic, $G$ is obtained from $H$ by adding an edge. If $H$ is $K_{4}$, clearly $G$ is $\theta$-based; if $H$ is $\mathbb{P}$, by Proposition 1.12, $G$ is $\theta$-based; in either case, we have a contradiction. This proves Theorem 1.27. ∎